OpenStudy (anonymous):

would you use quadratic formula, graphing, or factoring for 8x^2+117x-45=0 and how?

Directrix (directrix):

Factoring is always my first choice. That is, I look to see if the quadratic will factor over the set of Rational numbers. 8x^2+117x-45=0 (x + 15) (8x-3) = 0 ---> Factorization Zero Product Property x = -15 or 8x - 3 = 0 x = -15 or 8x = 3 and x = 3/8

OpenStudy (campbell_st):

this can be factorised multiply 8 and -45.... = -360... find the factors that add to 117 120 and - 3  gives the answer (x + 15)(8x - 3) x = -15 and 3/8 are solutions to the quadratic